A young entrepreneur is about to set up a new firm production hats. She calculates that the variable cost of producing each hat will be €2 and fixed costs of production will be €10.000. She will be selling her hats at 10€ each. How many units dose she need to sell to break even?
1
Expert's answer
2020-04-24T11:06:36-0400

"BEP=\frac{FC}{P-AVC}=\frac{10 000}{10-2}=1250"